I having trouble with this question because it is not following the regular D=r x t rule because of gravity. I can't find any thing else that helps me. Please let me know if I solved this correctly.
In a cartoon, a malfunctioning cannon fires a coyote towards the bottom of a cliff with an initial rate of 100 feet per second. If the cliff is 1250 feet tall, how long will it take the coyote to reach the desert floor? (To account for gravity, use the formula d=rt+16t squared)
1250=100t + 16t squared
t=6.25 seconds

Answer by stanbon(75887)

In a cartoon, a malfunctioning cannon fires a coyote towards the bottom of a cliff with an initial rate of 100 feet per second. If the cliff is 1250 feet tall, how long will it take the coyote to reach the desert floor? (To account for gravity, use the formula d=rt+16t squared)
1250=100t + 16t squared
t=6.25 seconds
------------------------------
The coefficient on t^2 takes care of the increase in speed due to gravity,
which is 16 ft/sec^2.
------------------------------
You only need to solve your equation for "t":
16t^2 +100t-1250 = 0
8t^2 + 50t - 625 = 0
t = [-50 +- sqrt(50^2 -4*8*-625)]/16
t = [-50 +- sqrt(22500)]/16
t = [-50- +- 150]/16
Positive answer:
t = 100/16 = 6.25 seconds
==================
Cheers,
Stan H.